A ladder is resting against the wall. The disatnce between the wall and the base of the ladder is 12 feet. What is the length of the ladder if the length is 6 feeet more than than the distance from the ground to the top of the ladder on the wall?
Answer by checkley77(12844)

Because this is a right triangle then:
a^2+b^2=c^2
12^2+x^2=(x+6)^2
144+x^2=x^2+12x+36
12x=144-36
12x=108
x^2=108/12
x=9 ft. ans. for the height of the ladder against the building.
(9+6)=15 ft. ans. for length of the ladder.
Proof:
12^2+9^2=15^2
144+81=225
225=225